# Credal Learning Theory

### v3

@@ 3.1.2 Belief functions as lower probabilities
A random set (Kendall, 1974; Matheron, 1975; Nguyen, 1978; Molchanov, 2005) is a set-valued random variable, modeling random experiments in which observations come in the form of sets. In the case of finite sample spaces, they are called belief functions (Shafer, 1976). While classical discrete mass functions assign normalized, non-negative values to the elements $\omega\in\Omega$ of their sample space, a belief function independently assigns normalized, non-negative mass values to subsets of the sample space: $m(A)\geq 0$, for all $A\subseteq\Omega$, $\sum_{A\subseteq\Omega}m(A)=1$. The belief function associated with a mass function $m$ then measures the total mass of the subsets of each event $A$,
$\text{Bel}(A)=\sum_{B\subseteq A}m(B).\vspace{-3pt}$ (1)
@@
$\vspace{-3pt}\mathcal{M}(\text{Bel})=\Big{\{}P:\Omega\rightarrow[0,1]:\text{Bel}(A)\leq P(A)\text{, }\forall A\subseteq\Omega\Big{\}},\vspace{-0pt}$ (2)
where $P$ is a data distribution. The dual upper probability to Bel is $\text{Pl}(A)\doteq 1-\text{Bel}(A^{c})$, for all $A\subseteq\Omega$. When restricted to singleton elements, it is called contour function, $\text{pl}(\omega)=\text{Pl}(\{\omega\})$.
@@ 3.1.3 Inferring belief functions from data
There are various ways one can infer a belief (or, equivalently, a plausibility) function from (partial) data, such as a sample of training sets. If a classical likelihood $\mathcal{L}$ having probability density or mass function (pdf/pmf) $\ell$ is available (as assumed in the frequentist paradigm), [^2] Here, pdf/pmf $\ell$ is the result of the Radon-Nikodym derivative of $\mathcal{L}$ with respect to a sigma-finite dominating measure $\nu$. one can build a belief function by using the normalized likelihood as its contour function. That is,
$\text{pl}(\omega)\doteq\frac{\ell(\omega)}{sup_{\omega^{\prime}\in\Omega}\ell(\omega^{\prime})},\quad\forall\omega\in\Omega,$ (3)
where $\Omega=\mathcal{X}\times\mathcal{Y}$ is the space where the training pairs live.
@@
As before, in our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we can specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$, and their corresponding pdf/pmf’s $\{\ell_{i}\}_{i=1}^{N}$. Then, we can compute ¯$\overline{\ell}(\omega)=\max_{i\in\{1,\ldots,N\}}\ell_{i}(\omega)$, for all $\omega\in\Omega$, and in turn $\text{pl}(\omega)=\overline{\ell}(\omega)/sup_{\omega^{\prime}\in\Omega}\overline{\ell}(\omega^{\prime})$, for all $\omega\in\Omega$ [^3] It is easy to see that pl is a well-defined plausibility contour. In turn, our credal set is derived as $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$, where $\text{d}P/\text{d}\nu=p$ is the pdf/pmf associated with distribution $P$ via its Radon-Nikodym derivative with respect to a sigma-finite dominating measure $\nu$. Such construction means that $\mathcal{P}$ includes all distributions whose pdf/pmf’s are element-wise dominated by plausibility contour pl.
3.1.4 Numerical example
@@
Let us give a simple numerical example. Let $\Omega=\{\omega_{1},\omega_{2},\omega_{3}\}$, where $\omega_{j}=(x_{j},y_{j})$, $j\in\{1,2,3\}$. Suppose also that we observed four training samples $D_{1},\ldots,D_{4}$ and that we specified the likelihood pmf’s $\ell_{1},\ldots,\ell_{4}$ as in Table 1.444In this case, $\nu$ is the counting measure. It is immediate to see that ¯$\overline{\ell}=(0.4,0.8,0.6)^{\top}$. [^5] We write the upper likelihood $\overline{\ell}$ in vector form for notational convenience. Then, by (3), we have that $\text{pl}=(0.5,1,0.75)^{\top}$. We then derive the lower and upper probabilities, $\underline{P}$ and $\overline{P}$ respectively, of $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$ on $2^{\Omega}$ as in Table 2 (Augustin et al., 2014, Section 4.4). As we can see, the probability bounds imposed by the credal set are not too stringent, and in line with the evidence encapsulated in $\ell_{1},\ldots,\ell_{4}$. Hence, the assumption that $P^{\text{true}}\in\mathcal{P}$ is very plausible. For a visual representation of $\mathcal{P}$, see the yellow convex region in Figure 2.
Figure 2: Credal set generated as in the example.
3.1.5 Fiducial approach
An alternative approach, proposed by Dempster and Almond (Almond, 1992), is based on fiducial inference (Hannig, 2009). Consider a parametric model, i.e. a family of conditional probability distributions of the data
$\Big{\{}f(\omega|\theta):\omega\in\Omega,\theta\in\Theta\Big{\}},$ (4)
where $\Omega$ is, again, the observation space and $\Theta$ is a parameter space. If the parametric (sampling) model (4) is supplemented by a suitably designed auxiliary equation $\omega=a(\theta,u)$, where $u$ is a “pivot” variable of known a-priori distribution $\mu$, one obtains a random set
$\Gamma(u)=\{(\omega,\theta)\in\Omega\times\Theta:\omega=a(\theta,u)\}\subseteq\Omega\times\Theta$ (5)
which, in turn, induces a belief function on the product space $\Omega\times\Theta$. This can be finally be marginalized to the data space $\Omega=\mathcal{X}\times\mathcal{Y}$ to generate a lower probability on it.
This approach was further extended by Martin, Zhang and Liu, who used a “predictive” random set to express uncertainty on the pivot variable itself, leading to an inference technique called weak belief (Zhang and Liu, 2011).
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we can derive $N$ many random sets $\Gamma_{i}$ as in (5), $i\in\{1,\ldots,N\}$, and consider the $N$ belief functions $\text{Bel}_{i}$ on $\Omega\times\Theta$ they induce. Then, we can compute their marginalization $\text{Bel}_{i}|_{\Omega}$ on the data space $\Omega=\mathcal{X}\times\mathcal{Y}$, and compute the minimum $\underline{\text{Bel}}|_{\Omega}\doteq\min_{i\in\{1,\ldots,N\}}\text{Bel}_{i}|_{\Omega}$. [^6] It is easy to see that $\underline{\text{Bel}}|_{\Omega}$ is itself a well-defined belief function In turn, our credal set is given by $\mathcal{P}=\mathcal{M}(\underline{\text{Bel}}|_{\Omega})$ as in equation (2).
@@ 3.2 Subjectivist Modeling
Another way of specifying a credal set is by taking a personalistic (or subjectivist) route (Walley, 1991; Caprio and Mukherjee, 2021). In this instance, the agent proceeds as follows. First, they specify the lower probability $\underline{P}_{\mathcal{S}}$ on a finite collection $\mathcal{S}=\{A_{\mathcal{S}}\}$ of subsets of $\mathcal{X}\times\mathcal{Y}$ – i.e. the smallest value that the probability of set $A_{\mathcal{S}}$ can take on, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$. This can be done, for example, as a result of the empirical distribution.
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, let $\mathcal{S}=\cup_{i=1}^{N}D_{i}$, and $P^{\text{emp}}_{i}$ denote the empirical distribution associated with $D_{i}$. Before going on, let us add a notational remark. At the beginning of the section, we denoted by $P_{i}^{\star}$ the true data generating process for training set $D_{i}$, $i\in\{1,\ldots,N\}$. The empirical distribution $P^{\text{emp}}_{i}$ is a (non-parametric) estimation of $P_{i}^{\star}$. On the other hand, recall that we denoted by $P^{\text{true}}\equiv P^{\star}_{N+1}$ the true data generating process for training set $D_{N+1}$. Fix any $A_{\mathcal{S}}\in 2^{\mathcal{S}}$; its (non-vacuous) lower probability $\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$ is defined as
$\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})\doteq\min\left\{{P^{\text{emp}}_{i}(A_{\mathcal{S}}):P^{\text{emp}}_{i}(A_{\mathcal{S}})>0}\right\}.$
The lower probability $\underline{P}_{\mathcal{S}}$ assigns the least nonzero value derived from the $N$ available empirical distributions, to any subset $A_{\mathcal{S}}$ of the whole available evidence $\mathcal{S}=\cup_{i=1}^{N}D_{i}$.
@@ 3.2.1 Walley’s Natural Extension
Once a lower probability $\underline{P}_{\mathcal{S}}$ on $2^{\mathcal{S}}$ is inferred, it can be (coherently uniquely) extended to a lower probability $\underline{P}$ on the whole sigma-algebra endowed to $\mathcal{X}\times\mathcal{Y}$ through an operator called natural extension (Walley, 1991, Sections 3.1.7-3.1.9), (Troffaes and de Cooman, 2014). The resulting extended lower probability is such that $\underline{P}(A_{\mathcal{S}})=\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$, and a lower probability value $\underline{P}(A)$ is assigned to all the other subsets $A$ of $\mathcal{X}\times\mathcal{Y}$ that are not in $\mathcal{S}$. It is also coherent – in Walley’s terminology – because, in the behavioral interpretation of probability derived from de Finetti (1974, 1975), its values cannot be used to construct a bet that would make the agent lose for sure, no matter the outcome of the bet itself.
Once $\underline{P}$ is obtained, the agent can consider the core of $\underline{P}$,
$\mathcal{M}(\underline{P})\doteq\{P:\underline{P}(A)\leq P(A)\text{, }\forall A\subseteq\mathcal{X}\times\mathcal{Y}\},$
i.e., the collection of all the (countably additive) probabilities that set-wise dominate $\underline{P}$. Marinacci and Montrucchio (2004); Cerreia-Vioglio et al. (2015) have shown that $\mathcal{M}(\underline{P})$ is indeed a (nonempty) credal set.
@@ 3.2.2 Properties of the Core
As shown in Amarante and Maccheroni (2006, Example 1) and Amarante et al. (2006, Examples 6, 7, 8), given a generic credal set $\mathcal{Q}$ whose lower envelope is $\underline{Q}$ – i.e. a credal set $\mathcal{Q}$ for which $\underline{Q}(\cdot)=\inf_{Q\in\mathcal{Q}}Q(\cdot)$ – we have that $\mathcal{M}(\underline{Q})\supseteq\mathcal{Q}$. From an information-theoretic perspective, this means that the uncertainty encapsulated in the core of a lower probability $\underline{Q}$ is larger than that in any credal set whose lower envelope is $\underline{Q}$ (Gong and Meng, 2021; Caprio and Mukherjee, 2023; Caprio and Gong, 2023; Caprio and Seidenfeld, 2023; Caprio et al., 2023b). In turn, $\mathcal{M}(\underline{Q})$ is the largest credal set the agent can build which represents their partial knowledge.
In our learning framework, given the available evidence $D_{1},\ldots,D_{N}$ that the agent uses to derive $\underline{P}_{\mathcal{S}}$, if the agent is confident that $P^{\text{true}}(A_{\mathcal{S}})\geq\underline{P}_{\mathcal{S}}(A_{\mathcal{S}})$, for all $A_{\mathcal{S}}\in 2^{\mathcal{S}}$, then it is natural to assume $P^{\text{true}}\in\mathcal{M}(\underline{P})$.
@@ 4 Generalization Bounds under Credal Uncertainty
Consider a credal set $\mathcal{P}$ on $\mathcal{X}\times\mathcal{Y}$ derived as in Section 3, and assume that we collect new evidence in the form of training set $D_{N+1}=\{(x_{N+1,1},y_{N+1,1}),\ldots,(x_{N+1,n_{N+1}},y_{N+1,n_{N+1}})\}$. To ease notation, in the following we refer to the newly acquired evidence as $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$.
An assumption that is common to all the results we present in this section is that $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P\equiv P^{\text{true}}$ i.i.d., and $P\in\mathcal{P}$. This means that either the new evidence comes from one of the distributions that generated $D_{1},\ldots,D_{N}$, or that it is compatible with the credal set we built from past evidence (Gong and Meng, 2021). That is, either $P$ set-wise dominates the lower probability $\underline{P}$ of $\mathcal{P}$ like in sections 3.1, 3.1.2, and 3.2, or it is set-wise dominated by the upper probability $\overline{P}$ (induced e.g. by a contour function) of $\mathcal{P}$ like in section 3.1.3. This is an extremely natural assumption, especially when the flow of training sets that we collect pertains to similar experiments or tasks. We note in passing that assuming $P\in\mathcal{P}$ is less stringent than what is typically done in frequentist statistics, where the data generating process is assumed to be perfectly captured by the likelihood. We instead posit that the true data generating process for the new evidence available belongs to a credal set $\mathcal{P}$, that was derived by the sample of training sets $D_{1},\ldots,D_{N}$. We also point out how it could be checked whether the assumption $P\in\mathcal{P}$ holds, similarly e.g. to what Cella and Martin (2022, Section 7) and Javanmardi et al. (2024) do for credal-set-based conformal prediction methods. That being said, PAC-like guarantees on the correct distribution $P$ being an element of the credal set $\mathcal{P}$ need to be developed, a task that we defer to future work.

### v4

@@ 3.1.2 Belief functions as lower probabilities
An alternative way to derive a credal set from the sample training evidence can be formulated within the framework of the Dempster-Shafer theory of evidence [27, 66].
A random set [39, 52, 55, 58] is a set-valued random variable, modeling random experiments in which observations come in the form of sets. In the case of finite sample spaces, they are called belief functions [66]. While classical discrete mass functions assign normalized, non-negative values to the elements $\omega\in\Omega$ of their sample space, a belief function independently assigns normalized, non-negative mass values to subsets of the sample space: $m(A)\geq 0$, for all $A\subseteq\Omega$, $\sum_{A\subseteq\Omega}m(A)=1$. The belief function associated with a mass function $m$ then measures the total mass of the subsets of each event $A$, $\text{Bel}(A)=\sum_{B\subseteq A}m(B)$.
@@
$\mathcal{M}(\text{Bel})=\{P:\Omega\rightarrow[0,1]:\text{Bel}(A)\leq P(A)\text{, }\forall A\subseteq\Omega\},$
where $P$ is a data distribution. The dual upper probability to Bel is $\text{Pl}(A)\doteq 1-\text{Bel}(A^{c})$, for all $A\subseteq\Omega$. When restricted to singleton elements, it is called the contour function, $\text{pl}(\omega)=\text{Pl}(\{\omega\})$.
@@ 3.1.3 Inferring belief functions from data
There are various ways one can infer a belief (or, equivalently, a plausibility) function from (partial) data, such as a sample of training sets. If a classical likelihood $\mathcal{L}$ having probability density or mass function (pdf/pmf) $\ell$ is available (as assumed in the frequentist paradigm), [^2] Here, pdf/pmf $\ell$ is the Radon-Nikodym derivative of $\mathcal{L}$ with respect to a sigma-finite dominating measure $\nu$. one can build a belief function by using the normalized likelihood as its contour function. That is, $\text{pl}(\omega)\doteq\frac{\ell(\omega)}{sup_{\omega^{\prime}\in\Omega}\ell(\omega^{\prime})}$, for all $\omega\in\Omega$, where $\Omega=\mathcal{X}\times\mathcal{Y}$ is the space where the training pairs live.
@@
As before, in our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we can specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$, and their corresponding pdf/pmf’s $\{\ell_{i}\}_{i=1}^{N}$. Then, we can compute ¯$\overline{\ell}(\omega)=\max_{i\in\{1,\ldots,N\}}\ell_{i}(\omega)$, for all $\omega\in\Omega$, and in turn
$\text{pl}(\omega)=\overline{\ell}(\omega)/sup_{\omega^{\prime}\in\Omega}\overline{\ell}(\omega^{\prime}),$ (1)
for all $\omega\in\Omega$ [^3] It is easy to see that pl is a well-defined plausibility contour function. In turn, our credal set is derived as $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$, where $\text{d}P/\text{d}\nu=p$ is the pdf/pmf associated with distribution $P$ via its Radon-Nikodym derivative with respect to a sigma-finite dominating measure $\nu$. Such construction means that $\mathcal{P}$ includes all distributions whose pdf/pmf’s are element-wise dominated by plausibility contour pl.
@@
Numerical example. Let $\Omega=\{\omega_{1},\omega_{2},\omega_{3}\}$, where $\omega_{j}=(x_{j},y_{j})$, $j\in\{1,2,3\}$. Suppose also that we observed four sample training sets $D_{1},\ldots,D_{4}$ and that we specified the likelihood pmf’s $\ell_{1},\ldots,\ell_{4}$ as in Table 1.444In this case, $\nu$ is the counting measure. There, we see e.g. how pmf $\ell_{1}$ assigns a probability of $0.3$ to the element $\omega_{1}$ of the state space $\omega$, and similarly for the other pmf’s and the other elements of the state space. It is immediate to see that ¯$\overline{\ell}=(0.4,0.8,0.6)^{\top}$. [^5] We write the upper likelihood $\overline{\ell}$ in vector form for notational convenience. denotes the transpose. Then, by Equation (1), we have that $\text{pl}=(0.5,1,0.75)^{\top}$.
We can then derive the lower $\underline{P}$ and upper $\overline{P}$ probabilities of $\mathcal{P}=\{P:\text{d}P/\text{d}\nu=p\leq{\text{pl}}\}$ on $2^{\Omega}$ as in Table 2, using the results in Augustin et al. [8, Section 4.4]. That is, $\underline{P}(A)=\max\left\{{\sum_{\omega\in A}\underline{P}(\omega),1-\sum_{\omega\in A^{c}}\overline{P}(\omega)}\right\}$ and $\overline{P}(A)=\min\left\{{\sum_{\omega\in A}\overline{P}(\omega),1-\sum_{\omega\in A^{c}}\underline{P}(\omega)}\right\}.$ As we can see from the visual representation of $\mathcal{P}$ (the yellow convex region in Figure 2), the probability bounds imposed by the credal set are not too stringent, and in line with the evidence encapsulated in $\ell_{1},\ldots,\ell_{4}$. Hence, the assumption that $P^{\text{true}}\equiv P^{\star}_{5}\in\mathcal{P}$ is quite plausible.
@@ 3.2 Subjectivist Modeling
Another way of specifying a credal set is by taking a personalistic (or subjectivist) route [14, 70]. In this approach, let $\{A_{\mathcal{S}}\}$ be a finite collection of subsets of $\Omega=\mathcal{X}\times\mathcal{Y}$. The agent first specifies the lower probability $\underline{P}_{\mathcal{S}}$ on the power set $2^{\mathcal{S}}$, where $\mathcal{S}=\cup A_{\mathcal{S}}$ – i.e., the smallest value that the probability of any subset of $\mathcal{S}$ can take on. This can be done, for example, as a result of the empirical distribution, as described below.
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, we have that $\{A_{\mathcal{S}}\}=\{D_{i}\}_{i=1}^{N}$, and so $\mathcal{S}=\cup_{i=1}^{N}D_{i}$. Recall that we originally denoted by $P_{i}^{\star}$ the true data generating process for training set $D_{i}$, $i\in\{1,\ldots,N\}$: the empirical distribution $P^{\text{emp}}_{i}$ is a (non-parametric) estimation of $P_{i}^{\star}$. On the other hand, recall that we denoted by $P^{\text{true}}\equiv P^{\star}_{N+1}$ the true data generating process for the test set of data $D_{N+1}$.
The lower probability $\underline{P}_{\mathcal{S}}$ is defined as follows. For every element $(x,y)$ in $\mathcal{S}=\cup_{i=1}^{N}D_{i}$, we let $\underline{P}_{\mathcal{S}}(\{(x,y)\})\doteq\min\{P^{\text{emp}}_{i}(\{(x,y)\}):P^{\text{emp}}_{i}(\{(x,y)\})>0\}$. Requiring $\underline{P}_{\mathcal{S}}(\{(x,y)\})=\min_{i}P^{\text{emp}}_{i}(\{(x,y)\})$ is not enough because, if the training data sets do not overlap, we would end up having lower probability 00 for some singleton that we observed at training time, and hence we would be neglecting some collected evidence. The lower probability $\underline{P}_{\mathcal{S}}$ of all the other non-singleton elements $B$ of $\mathcal{S}$ is computed according to [8, Equation (4.6a)], that is,
$\underline{P}_{\mathcal{S}}(B)=\max\left\{{\sum_{(x,y)\in B}\underline{P}_{\mathcal{S}}(\{(x,y)\}),1-\sum_{(x,y)\in B^{c}}\max_{i}P^{\text{emp}}_{i}(\{(x,y)\})}\right\}.$ (2)
Numerical example. Suppose for simplicity that $\mathcal{X}=\{x\}$, so that $\Omega=\{x\}\times\mathcal{Y}\simeq\mathcal{Y}$, and let $\mathcal{Y}=\{1,\ldots,10\}$. Suppose $N=2$, and let $D_{1}$ be a collection of three $4$’s, three $5$’s, and three $6$’s. Let also $D_{2}$ be a collection of six $5$’s and two $6$’s. Then, $\mathcal{S}=\{4,5,6\}$ and $2^{\mathcal{S}}=\{\emptyset,\{4\},\{5\},\{6\},\{4,5\},\{5,6\},\{4,6\},\{4,5,6\}\}$. In turn, $\underline{P}_{\mathcal{S}}(\{4\})=1/3$, $\underline{P}_{\mathcal{S}}(\{5\})=1/3$, and $\underline{P}_{\mathcal{S}}(\{6\})=1/4$. By (2), this implies that $\underline{P}_{\mathcal{S}}(\{4,5\})=2/3$, $\underline{P}_{\mathcal{S}}(\{5,6\})=2/3$, and $\underline{P}_{\mathcal{S}}(\{4,6\})=\max\{1/3+1/4,1-\max_{i\in\{1,2\}}P^{\text{emp}}_{i}(\{5\})\}=\max\{7/12,1-\max\{1/3,3/4\}=\max\{7/12,1-3/4\}=7/12$. Of course, $\underline{P}_{\mathcal{S}}(\emptyset)=0$ and $\underline{P}_{\mathcal{S}}(\mathcal{S})=1$.
@@ 3.2.1 Walley’s Natural Extension
Once a lower probability $\underline{P}_{\mathcal{S}}$ on $2^{\mathcal{S}}$ is inferred, it can be (coherently uniquely) extended to a lower probability $\underline{P}$ on the whole sigma-algebra endowed to $\mathcal{X}\times\mathcal{Y}$ through an operator called natural extension [69], [70, Sections 3.1.7-3.1.9]. The resulting extended lower probability is such that $\underline{P}(B)=\underline{P}_{\mathcal{S}}(B)$, for all $B\in 2^{\mathcal{S}}$, and a lower probability value $\underline{P}(A)$ is assigned to all the other subsets $A$ of $\mathcal{X}\times\mathcal{Y}$ that are not in $\mathcal{S}$. It is also coherent – in Walley’s terminology – because, in the behavioral interpretation of probability derived from de Finetti [25, 26], its values cannot be used to construct a bet that would make the agent lose for sure, no matter the outcome of the bet itself.
[…]
@@ 3.2.2 Properties of the Core
As shown in Amarante and Maccheroni [5, Example 1] and Amarante et al. [6, Examples 6, 7, 8], given a generic credal set $\mathcal{Q}$ whose lower envelope is $\underline{Q}$ – i.e., a credal set $\mathcal{Q}$ for which $\underline{Q}(\cdot)=\inf_{Q\in\mathcal{Q}}Q(\cdot)$ – we have that $\mathcal{M}(\underline{Q})\supseteq\mathcal{Q}$. From an information-theoretic perspective, this means that the uncertainty encapsulated in the core of a lower probability $\underline{Q}$ is larger than that in any credal set whose lower envelope is $\underline{Q}$ [13, 15, 16, 18, 30]. In turn, $\mathcal{M}(\underline{Q})$ is the largest credal set the agent can build which represents their partial knowledge. In our learning framework, given the available evidence $D_{1},\ldots,D_{N}$ that the agent uses to derive $\underline{P}_{\mathcal{S}}$, if the agent is confident that $P^{\text{true}}(B)\geq\underline{P}_{\mathcal{S}}(B)$, for all $B\in 2^{\mathcal{S}}$, then it is natural to assume $P^{\text{true}}\equiv P^{\star}_{N+1}\in\mathcal{M}(\underline{P})$.
@@ 4 Generalization Bounds under Credal Uncertainty
Consider a credal set $\mathcal{P}$ on $\mathcal{X}\times\mathcal{Y}$ derived as in Section 3, and assume that we collect new evidence in the form of a test set of data $D_{N+1}=\{(x_{N+1,1},y_{N+1,1}),\ldots,(x_{N+1,n_{N+1}},y_{N+1,n_{N+1}})\}$. To ease notation, in the following we refer to the newly acquired evidence as $(x_{1},y_{1}),\ldots,(x_{n},y_{n})$.
An assumption that is common to all the results we present in this section is that $(x_{1},y_{1}),\ldots,(x_{n},y_{n})\sim P\equiv P^{\text{true}}\equiv P^{\star}_{N+1}$ i.i.d., and $P\in\mathcal{P}$. This means that either the new evidence comes from one of the distributions that generated $D_{1},\ldots,D_{N}$, or that it is at least compatible with the credal set we built from past evidence [30]. That is, either $P$ set-wise dominates the lower probability $\underline{P}$ of $\mathcal{P}$ (as in Sections 3.1, 3.1.2, and 3.2), or it is set-wise dominated by the upper probability $\overline{P}$ of $\mathcal{P}$ (induced, e.g., by a contour function as in Section 3.1.3). This is a rather natural assumption, especially when the stream of training sets we collect pertains to similar experiments or tasks. For example, this is the case in Continual Learning (CL), where it is customary to assume task similarity [47], that is, to posit that the oracle distributions pertaining to all the tasks of interest are all contained in a TV-ball of radius chosen by the user. A more complete discussion on the relation between our credal approach and CL can be found in Appendix F. It is also the case in the healthcare setting, where experts’ opinions can be incorporated alongside empirical data (plausible probability distributions) to represent the probability uncertainty, for example, for the prognosis of a disease given a set of patient characteristics/biomarkers [65]. To make sure that the credal set constructed encapsulates most of the “potential distributions needed”, a number of approaches can be taken including incremental learning; in this approach the AI model learns and updates knowledge incrementally. As a result, the credal set can be continuously updated (via incremental learning) as new data become available. In this direction, learning health systems are being implemented in practice. These are health systems “in which internal data and experience are systematically integrated with external evidence, and that knowledge is put into practice”.
We note, in passing, that assuming $P\in\mathcal{P}$ is less stringent than what is typically done in frequentist statistics, where the data-generating process is assumed to be perfectly captured by the likelihood. We instead posit that the true data generating process for the new evidence available belongs to a credal set $\mathcal{P}$, that was derived by the sample of training sets $D_{1},\ldots,D_{N}$.
Formal ways of checking whether the assumption $P\in\mathcal{P}$ holds exist, e.g., by following what Cella and Martin [19, Section 7] and Javanmardi et al. [36] do for credal-set-based conformal prediction methods, or more general approaches [1, 22, 29, 46, 56]. [^6] If we want to avoid to formally check the assumption that $P\in\mathcal{P}$, we need to show on a case-by-case basis that either the credal set covers a non-negligible portion of the distribution class of interest, or that even a small credal set is “good enough” for the analysis at hand. That being said, deriving PAC-like guarantees on the correct distribution $P$ being an element of the credal set $\mathcal{P}$ is a desirable objective - a task that we defer to future work. Here, we focus on formally deriving what the consequences are in terms of generalization bounds.